Adult tickets for a play cost $5 and children tickest cost $2 for 1person. 460 tickets were sold. Recipts for person were $1880. Find the # of adult tcikets sold. You must use the chart.(first colmn in chart is type, than cost,and lastly value) the answer is 320. But I dont know what they plugged into the chart to get this answer. I have the types as Adult, children, and mixture. Cost as 5,2,and 7. And the value as 5x, 2(460-%x) and (460)7. What am I doing wrong?
Answer by ptaylor(1894)

Let x=the number of adult tickets sold
Then 460-x=the number of children tickets sold
We know that the amount of $ collected from selling adult tickets (x)($5) plus the amt of $ collected from selling children tickets (460-x)($2) equals $1880. The equation to solve is:
5x+(460-x)(2)=1880 simplifying, we get
5x-2x=1880-920
3x=960
x= 320 adult tickets
I think you may be over complicating the problem. You don't need to worry about the mixture. That takes care of itself.
USING THE CHART, WE HAVE:
Column I (Type) will have two entries: x(Number of adult tickets) and 460-x (Number of children tickets); Column II(Cost) will have two entries: $5 (Cost of adult tickets) and $2 (Cost of children tickets); Column III (Value) will also have two entries 5x (value of adult tickets sold) and (460-x)(2)(value of children tickets sold). Adding Column III to find the total value which will equal 1880, we have:
5x+(460-x)(2)=1880 The same equation as above
Hope this helps----ptaylor